OpenStudy (anonymous):

What is the y-intercept of the equation 2x-5y=-30? A.-15 B.-6 C.6 D.15

Directrix (directrix):

B.-6 because...

OpenStudy (katrinakaif):

First Turn the equation in y=mx + b format 2x - 5y = -30 -5y = -30 + 2x y = -30/-5 + 2x/-5 y = 2/5x + 6 y = mx + b B = y Intercept. C

OpenStudy (anonymous):

Thank you KatrinaKaif.

OpenStudy (katrinakaif):

The pleasure is mine.

OpenStudy (anonymous):

Directrix, thank you....for I'm not sure yet.

Directrix (directrix):

What is the y-intercept of the equation 2x-5y=-30? 2x - 5y = -30 -5y = -30 - 2x y = -30/(-5) - (2x/(-5) y = 6 - (2/5)x y-intercept is 6

Directrix (directrix):

You are welcome, Tsterzz. I am sure of that.